If X is a random variable such that E(X)=3 and V(X)=2, and if Y is a random variable such that Y=6+2X. Calculate the mean and variance of Y.

a) E(Y)=12

b) V(Y)=
Answer #1

we want to find V(Y)

given value  E(X)=3   V(X)=2

Y=6+2X.

V(Y)=V(6+2*x) = 2^2 * V(x) = 4*2 = 8 ( we know the constant of variance is zero)